pablo:

Two studies were completed in California. One study in northern California involved 1,000 patients; 74% of them experienced flulike symptoms during the month of December. The other study, in southern California, involved 500 patients; 34% of them experienced flulike symptoms during the same month. Which study has the smallest margin of error for a 98% confidence interval?

Vocaloid:

Moved to Statistics

sillybilly123:

According to the Central Limit Theorem, sample proportion, \(\hat p\), is distributed normally as: \(~~~ \hat p \sim \mathbb N\left(\mu_{\hat p} = p, \sigma_{\hat p} = \sqrt{\dfrac{p(1-p)}{n}} \right) \) ... where \(n\) is the related sample size The MoE is: \(~~~~~ z^{*} \cdot \sqrt{\dfrac{\hat p ( 1 - \hat p)}{n}}\) For \(98 \%\) confidence level, \(z^{*} = 2.33\) - from the standard normal distribution table For the first sample, MoE is: \( ~~~~~ 2.33 \cdot \sqrt{\dfrac{0.74 ( 1 - 0.74)}{1000}} = 3.2 \% ~ ~ [1dp]\) (ie TBC, we conclude, with \(98 \%\) confidence, that \(74\% \pm 3.2 \%\) of that **population** will have experienced flu-like symptoms in Dec) For the second sample, MoE is: \(~~~~~ 2.33 \cdot \sqrt{\dfrac{0.34 ( 1 - 0.34)}{500}} = 4.9 \% ~ ~ [1dp]\) ---> The first study has the smallest MoE for **any** given confidence level. Sample size is key .... more people you test, more likely you are to get to truth.